### v1

# InterActHuman: Multi-Concept Human Animation with Layout-Aligned Audio Conditions

Abstract
End-to-end human animation with rich multi-modal conditions, e.g., text, image and audio has achieved remarkable advancements in recent years. However, most existing methods could only animate a single subject and inject conditions in a global manner, ignoring scenarios that multiple concepts could appears in the same video with rich human-human interactions and human-object interactions. Such global assumption prevents precise and per-identity control of multiple concepts including humans and objects, therefore hinders applications. In this work, we discard the single-entity assumption and introduce a novel framework that enforces strong, region-specific binding of conditions from modalities to each identity’s spatiotemporal footprint. Given reference images of multiple concepts, our method could automatically infer layout information by leveraging a mask predictor to match appearance cues between the denoised video and each reference appearance. Furthermore, we inject local audio condition into its corresponding region to ensure layout-aligned modality matching in a iterative manner. This design enables the high-quality generation of controllable multi-concept human-centric videos. Empirical results and ablation studies validate the effectiveness of our explicit layout control for multi-modal conditions compared to implicit counterparts and other existing methods. See our Project Page for more details.
1 Introduction
By leveraging the priors of pretrained Diffusion Transformer-based (DiT) video diffusion models bar2024lumiere ; svd ; ayl ; guo2024animatediff ; zhou2022magicvideo ; walt ; wang2023modelscope ; vdm ; videogan ; cvideogan ; singer2022make ; text2video ; villegas2022phenaki ; lin2025apt , end-to-end human animation models, especially audio-driven approaches he2023gaia ; tian2024emo ; xu2024hallo ; wang2024vexpress ; chen2024echomimic ; xu2024vasa ; stypulkowski2024diffused ; jiang2024loopy ; lin2024cyberhost ; lin2025omnihuman has achieved high-quality human-centric video generation and strong controllability from multi-modal conditions, such as text, image and audio. However, most of existing methods commonly hold an assumption of a single-identity paradigm: all available conditions should be fused globally and implicitly assumed to describe one unique subject in the given image. Although this global injection strategy simplifies conditioning by sharing the same condition signal across all regions, it fundamentally limits scalability in scenarios involving multiple individuals or complex human-object interactions, where each entity requires distinct appearance and voice attributes.
Figure 1: Video frames generated from audio and multi-concept reference images (human heads/full bodies, objects, scenes) display rich, audio-matched expressions. Our method enables compositional generation including outfit changes, human–object interactions, anime styles, dialogues even without a start frame. Red and green wave icons denote speaking and listening, respectively.
Recent multi-concept video customization methods, such as Video-Alchemist chen2025multi , ConceptMaster huang2025conceptmaster , Phantom liu2025phantom , and SkyReels-A2 fei2025skyreels , enable injecting multiple reference images into a single video, facilitating multi-person or human-object interaction scenarios. However, these methods face significant challenges when directly applied to human animation tasks. In such tasks, animation signals are highly specific to individual identities and demand precise conditioning and alignment with specific spatiotemporal regions. For example, audio signals are exclusively associated with the current speaker and are unrelated to the background or other individuals’ concepts. Existing customization methods, akin to end-to-end human animation approaches, still adopt video-level condition injection. While this approach may suffice for general video generation, it often causes confusion in human-centric video generation, making it difficult to produce satisfactory results, as shown in Fig. 3. For end-to-end human-centric video generation, conditioning inputs should ideally include not only global modalities (e.g., reference images and text descriptions) but also local modalities (e.g., audio). As mentioned, existing human animation and multi-concept customization methods fail to address this critical distinction. This limitation motivates us to propose a new framework that enables the precise injection of local human-related modalities, a capability that is both essential and urgently needed for robust multi-concept, human-centric video generation, as illustrated in Fig. 1.
In this paper, we propose InterActHuman, a video diffusion framework to achieve spatial alignment of multi-modal conditions for multi-concept human video generation. Unlike previous methods chen2025multi ; huang2025conceptmaster , which rely on feature fusion and attention mechanisms to implicitly learn the relationships between condition signals and concepts, InterActHuman introduces an attention module that explicitly predicts the spatial locations where reference concepts appear in the video. This explicit alignment allows the model to accurately associate local audio conditions with the correct spatial regions via iterative mask prediction and masked audio attention during the diffusion inference process. Our framework offers two key advantages. First, it enforces a stronger layout constraint by explicitly and precisely binding conditions to their corresponding spatial regions. Second, it provides a unified interface for synchronously injecting all modalities such as visual and acoustic inputs, through the layout. These features make InterActHuman well-suited for multi-modal multi-concept human animation and serve as a baseline framework for this domain.
Although using an explicit mask for condition injection might seem straightforward at first glance, it introduces an intrinsic chicken-and-egg dilemma: during inference, the final video is not yet available, leaving the spatial positions of each identity uncertain, thus making accurate mask prediction impossible; yet without those masks, the spatial injection of local audio conditions cannot be achieved, resulting in an incomplete or misaligned video generation process. To address this dilemma, we leverage the iterative denoising process inherent in diffusion models song2021ddim . Specifically, we introduce a mask predictor branch in the diffusion pipeline and adopt an interleaved mask prediction strategy, wherein the mask predictions generated at step $k$ serve as guidance for condition injection in step $k+1$. This iterative refinement breaks the cyclic dependency by progressively ‘discovering’ the spatial layout, ensuring that even in the absence of ground-truth video during inference, precise spatial alignment is achieved. In essence, our approach transforms the chicken-and-egg problem into a sequential and converging process that robustly aligns local audio conditions.
Beyond model design, we developed a scalable pipeline to automatically assemble high-quality, human-centric animation data to address the lack of suitable multi-concept datasets. This pipeline operates by: 1) accurately tracking individual identities to extract their mask information and images, and 2) aligning audio segments to each identity through lip synchronization. We curated a dataset of over two million video-entity pairs, capturing both human-human and human-object interactions across a wide range of object categories. In summary, our contributions are as follows:
(1) We propose a novel human animation framework capable of synthesizing multi-person and human-object interactions, conditioned on multiple reference images, text descriptions, and audio inputs. The framework also supports long video generation as well as single full-image conditioning. (2) We highlight the importance of local condition injection for multi-concept, multi-modal video generation and introduce a simple yet effective design that enables the model to handle both global and local conditions by automatically localizing the conditioned layout. Experimental results demonstrate that our proposed design significantly outperforms existing baselines.
2 Related Works
Video Diffusion Models have enabled unprecedented quality text-to-video (T2V) or image-to-video (I2V) generation in recent years, thanks to diffusion-based generative models jonathan2020ddpm ; song2021ddim ; karras2022edm ; song2020score ; liu2022flow . Early T2V approaches either adapt pretrained text-to-image networks in a training-free manner singer2022make ; wu2023tune ; qi2023fatezero or fine-tune UNet-based latent diffusion architectures guo2024animatediff ; zhou2022magicvideo ; svd ; wang2023modelscope . To push the frontier, recent works compress spatiotemporal features with 3D causal VAEs yu20233DVAE , and migrate to Diffusion Transformer (DiT) backbones vaswani2017attention ; peebles2023scalable ; videoworldsimulators2024 ; hong2022cogvideo . Through progressive low-to-high resolution pretraining and fine-tuning polyak2024moviegen ; kong2024hunyuanvideo ; wang2025wan , these models yield longer, more coherent, and high-quality videos. Our method is also built upon the pretrained DiT video generation models.
Human Animation Models synthesize videos of people driven by text, reference images, human body poses or audios. Early GAN-based methods siarohin2019fomm ; zhao2022tps ; siarohin2021mraa ; jiang2024mobileportrait ; wang2021facev2v are trained on small datasets nagrani2017voxceleb ; siarohin2019fomm ; xie2022vfhq ; zhu2022celebv for self-supervised pose transfer. Diffusion-based approaches shao2024human4dit ; zhang2024mimicmotion ; aa ; DBLP:conf/nips/00010ZF0LTCX0L24 ; wang2025multi now surpass GAN-based methods by conditioning on 2D skeletons, 3D depth, or mesh sequences. Audio-driven portrait methods GeneFace ; zhang2023sadtalker ; EMO ; jiang2024loopy ; hallo3 have been extended toward full-body motion via two-stage pipelines for improved hand quality VLogger ; EchomimicV2 ; EMO2 ; diffted and one-stage unified framework designs lin2024cyberhost ; lin2025omnihuman . In summary, none of them explored multi-concept human animation, and our method is the first to enable multi-concept human animation with local audio conditions.
Multi-Concept Video Customization Models have received limited attention. Early single-concept identity-preserving methods include Videobooth jiang2024videobooth , which learns coarse-to-fine embeddings from WebVid bain2021frozen via Grounded-SAM kirillov2023segment ; liu2023grounding ; ren2024grounded ; ID-Animator he2024id , which integrates IP-Adapter ye2023ip into AnimateDiff guo2024animatediff ; and ConsisID yuan2024identity , which decouples frequency signals to preserve facial identity. More recently, Video-Alchemist chen2025multi , ConceptMaster huang2025conceptmaster , and Phantom liu2025phantom support multiple reference images and text descriptions via cross- or self-attention injection for general-purpose multi-concept customization. BlobGen-Vid feng2025blobgen further enables local layout control of text-specified concepts using user- or LLM-provided spatiotemporal masks. However, all existing methods rely solely on image and text conditions and lack support for multimodal inputs such as audio, which we argue are essential for truly versatile, human-centric video generation.
3 Method
In this section, we present InterActHuman, our multi-concept video generation framework designed to address the challenges of local condition matching for identities in multi-modal conditions. As shown in Fig. 2, the framework begins by customizing reference images of multiple concepts into a video and predicting the mask regions corresponding to each reference appearance in the output video as layout cues with a cross-attention as mask predictor. It operates on the features of the noisy video latent and the reference image latent and supervised by ground-truth masks. Finally, local audio conditions are injected into the designated regions guided by an iterative mask prediction procedure.
3.1 Problem Setting and Preliminaries
Problem Setting. Given a caption $T$ that describes a target video, alongside a set of concept reference images $\{X_{i}\mid i=1,\ldots,N\}$ and identity-level audio $\{Y_{i}\mid i=1,\ldots,N\}$ where $N$ denotes the number of distinct concepts, the objective of Multi-Concept Human Animation is to generate high-quality videos that faithfully integrate all image-specified visual concepts with correct lip movements synchronized with each audio signal in accordance with the descriptive caption $T$. Each concept should consistently preserve its visual identity as depicted in the provided images, while accurately expressing the semantic roles and behaviors described in the caption.
Preliminaries. Transformer-based text-to-video latent diffusion models have recently demonstrated remarkable capabilities in generating high-quality video content. Our proposed InterActHuman framework is built upon the MMDiT-based video generation model dit ; esser2024sd3 ; seawead2025seaweed and utilize a 3D Variational Autoencoder (VAE) kingma2013auto , which compresses input videos into a compact latent space. For training, we adopt the flow matching objective lipman2022flow , which formulates the generative process as a probability flow ordinary differential equation (ODE). This ODE transports clean latent representations $z_{0}$ to their noisy counterparts $z_{t}$ along a linear path, defined as $z_{t}=(1-t)z_{0}+t\epsilon$ at timestep $t$, where $\epsilon$ is sampled from a standard Gaussian distribution. In our setting, which incorporates multi-modal conditions (image and audio), the output of the diffusion transformer is parameterized as $v_{\Theta}(z_{t},t,c_{img},c_{audio})$. This output is supervised to predict the velocity $(z_{1}-z_{0})$, resulting in the following training objective:
$\mathcal{L}=\mathbb{E}_{t,z_{0},\epsilon}\left||v_{\Theta}(z_{t},t,c_{img},c_{audio})-(z_{1}-z_{0})\right||_{2}^{2}.$
This formulation ensures the model effectively learns to generate videos conditioned on multi-modal inputs while maintaining high fidelity and temporal consistency.
Figure 2: Illustration of our framework, which adaptively predicts masks as the spatial guidance of audio condition injection. In training, we train the mask predictor (cross-attn w/ MLP) with mask loss; in inference, we collect mask predictions to cache and leverage masks predicted from the last denoising step ($t-1$) to guide the audio cross-attn in the current denoising step ($t$).
3.2 Multi-Concept Reference Image Injection
To handle multiple appearance images, we inject appearance cues via self-attention in the original DiT layers. Each reference image $X_{i}$ is encoded into latent tensors ${\mathbf{x}_{i}}$ using the same VAE as for noisy video frames. The reference latents are then flattened into token sequences and processed by the DiT along with the noisy latents $\mathbf{v}$. The reference latents will reuse the parameters of DiT to extract features and interact with noisy latents at the self-attention layer in every DiT block, allowing appearance cues to propagate to the denoising pathway. Notably, no extra networks or additional parameters are required, thus preserving the model’s efficiency.
3.3 Layout Prediction and Local Condition Injection
Although reference image injection has been addressed, leveraging spatiotemporal layouts for local conditions like precise audio alignment remains challenging. There is a conflict between mask prediction and audio injection: if the forward pass is incomplete, masks are unavailable; if it is complete, it’s too late for local condition injection. We resolve this by using the multi-step inference process of diffusion models song2021ddim , where masks predicted at step $k$ guide multi-modal condition injection at step $k+1$. As our network learns spatiotemporal layouts for each identity, we obtain precise masks without requiring user annotations.
Mask Predictor. Within each DiT block, we predict a spatio-temporal mask that quantifies how strongly each reference image should influence each video frame. We attach a lightweight mask-predictor head to each of the $L$ transformer layers. The head comprises: (1) a shared linear projection that maps the hidden video features $\mathbf{h}^{v}$ and the hidden reference features $\mathbf{h}^{r}_{i}$ to query, key, and value tensors, (2) a LayerNorm, (3) 3D RoPE positional encoding, (4) a cross-attention module, and (5) a two-layer MLP. After normalization and RoPE, we let the video tokens attend to a single reference by computing $a_{i}^{(l)}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{v}{\mathbf{K}^{r}_{i}}^{\top}}{\sqrt{d}}\right)\mathbf{V}^{r}_{i}$, where $d$ is the head dimension and $l$ indexes the current layer. The attended feature $\mathbf{a}^{(l)}_{i}$ is subsequently transformed by the MLP and passed through a sigmoid to yield a layer-specific mask $\mathbf{m}^{(l)}_{i}\in[0,1]^{T}$ for reference $X_{i}$. Then we average the mask prediction from last few DiT blocks to form the final mask. Notably, the mask predictor is always trained to predict complete region of a human, regardless of whether the reference image contains only the upper body, face, or full body of a person. This design simplifies the mask prediction task and improves consistency for the conditioning module, ensuring robust and reliable conditioning across diverse scenarios. Moreover, since the mask predictors reuse in-block features and share parameters across references, they introduce minimal computational overhead while enabling explicit spatiotemporal layout control for each identity.
Mask Prediction by Caching During Inference.During inference, mask prediction requires the integration of results across all layers, which poses challenges in the early stages of denoising. At these initial steps, the latent representations of reference concepts are often too ambiguous, preventing the conditioning module from generating precise mask layouts. To address this issue, we adopt an iterative mask prediction strategy. At each denoising step, the mask predicted in the previous step is cached as a layout guide for the current step, enabling the model to progressively improve the accuracy of spatiotemporal localization. This iterative process ensures that the masks dynamically adapt to the evolving video content, maintaining consistency for each concept throughout the sequence and improving the overall conditioning quality.
Local Audio Conditioning. To prepare our model for multi-concept human-centric video generation, we first pretrain it on single-identity audio-conditioned animation by adding cross-attention-based audio conditioning without masks and a mixed-conditions training strategy, following recent work lin2025omnihuman . It utilizes a new cross-attention layer to inject wav2vec baevski2020wav2vec2 audio features into each DiT block after the MMDiT layer. In multi-concept setting, we find that utilizing pretrained audio cross-attention has already lead to satisfactory per-identity audio-driven performance. So we implement local audio conditioning mainly in the iterative process in inference. Instead of processing the entire set of noisy video tokens as in previous works lin2025omnihuman ; lin2024cyberhost , we inject wav2vec features into the tokens corresponding to a specific identity, using the mask predicted in the previous step as guidance. To ensure smooth transitions in noisy video features and final output videos, we apply a linear combination of meaningful audio features and muted audio features for each noisy video token, based on the predicted mask confidence along the mask edges.
Moreover, such local audio conditioning enables us to generate multi-person videos with dialogues. As long as the dialogue audio are separated into individual audios as the input, our model could generate realistic multi-person videos where people take turns to speak, facilitating many applications.
Training Loss and Strategies. The overall training loss combines a flow matching diffusion loss (see Sec. 3.1) with a focal losslin2017focal for mask classification, which mitigates the training instability and gradient explosion encountered with binary cross-entropy loss, likely due to foreground-background imbalance or low-quality masks. A frame alignment flag is employed to exclude frames with invalid or low-quality masks from loss computation, ensuring robust temporal supervision. To address the ‘copy-paste behavior’ of diffusion models, which often replicate subjects from reference images with minimal variation in camera angle or posture, we apply random masking to reference images, selectively revealing the head, full body, or clothing of the original subject appearance to encourage appearance diversity. We use an appearance conditioning ratio of face to full body of $0.7\!:\!0.3$, while balancing the contributions of the diffusion loss and focal loss equally with a ratio of $1\!:\!1$.
Inference Strategies. Given $N$ reference images and a text prompt, our method use Qwen2.5-VL bai2025qwen2 as the re-phraser to get detailed descriptions from each reference images, and then integrate them with the original text prompt to form the final prompt. We utilize a shared classifier-free guidance (CFG) across audio and text. To achieve local control, audio conditions are injected solely within the predicted mask regions for each reference. Notably, early denoising steps yield unreliable masks, and using them for condition injection can suppress other references. Thus, mask-based condition injection is omitted for the first 10 denoising steps and applied subsequently using the mask predicted from the previous step. We only inject audio signals with masks into positive result while applying CFG, and use a CFG strength as 6.5 and 50 denoising steps.
Figure 3: Qualitative comparison with previous methods on multi-concept audio injection.
3.4 Multi-Concept Human-Centric Video Data Curation
We curate a large-scale, human-centric video dataset from a large internal video dataset, filtering out videos that are too short (<4 seconds) or contain too many salient humans (>3) as detected by a human pose detector dwpose . Unlike previous multi-concept video generation methods chen2025multi ; huang2025conceptmaster that rely on conventional pipelines like Grounding-DINO liu2023grounding and SAM kirillov2023segment for obtaining a single reference image per identity, our multi-stage pipeline leverages advanced vision-language models. First, each raw video is densely captioned using Qwen2-VL wang2024qwen2 (distilled from Gemini-2.0-Pro), generating fine-grained descriptions of the environment, subjects’ appearance, actions, expressions, interacting salient objects, and inter-subject interactions. Next, these descriptions are parsed by a zero-shot Gemini-2.0-Flash API to extract structured appearance phrases. For spatial supervision, we use Grounding-SAM2 ren2024grounded with the query person to produce accurate, temporally consistent masks, which are used both for extracting foreground reference images (with a white background) and as ground-truth for our mask predictor. Our corpus comprises over $2.6$M triplets of videos, per-frame masks, and captions, forming the foundation of InterActHuman.
4 Experiments
Baselines. Since ConceptMaster huang2025conceptmaster and Video-Alchemist chen2025multi are not open-sourced, we do not have access to their models nor test sets. Therefore, we primarily compare our method with recent multi-concept video customization models through their publicly available APIs or public models, evaluating performance from the perspectives of visual appearance and adherence to text prompts, including Vidu2.0 bao2024vidu , Pika2.1 pikaart2024 , Kling 1.6 kuaishou2024 and Phantom liu2025phantom , and audio-driven human video generation methods, including DiffTED diffted , DiffGest zhu2023taming + Mimiction zhang2024mimicmotion , CyberHost lin2024cyberhost , OmniHuman lin2025omnihuman and Kling 1.6 kuaishou2024 with lip synchronization.
Table 1: Quantitative comparisons with audio-conditioned full-body animation baselines.
Methods Single-Person Test Set Multi-Person Test Set
[…]
DiffTED diffted 0.926 - 0.769 - - - -
DiffGest. zhu2023taming +Mimic. zhang2024mimicmotion 0.496 23.409 0.833 - - - -
CyberHost lin2024cyberhost 6.627 24.733 0.884 8.974 4.011 2.856 54.797
Kling1.6 kuaishou2024 + Lip-sync. 4.449 46.490 0.826 8.401 4.716 3.444 33.555
OmniHuman lin2025omnihuman w/o mask 7.443 47.561 0.898 9.482 4.768 3.466 33.895
OmniHuman lin2025omnihuman w/ fixed mask - - - 7.068 4.690 3.369 40.239
Ours 7.272 59.635 0.885 6.670 4.757 3.467 22.881
Table 2: User preference evaluation. means publicly available version with Wan2.1-1.3B.
Audio-Driven Multi-Concept Customization
Metric Kling kuaishou2024 OmniHuman lin2025omnihuman Ours Pika pikaart2024 Phantom liu2025phantom Kling kuaishou2024 Vidu bao2024vidu Ours
Avg. Score $\uparrow$ 1.70 1.82 2.48 2.22 2.46 2.90 3.40 4.01
Top-1 (%) $\uparrow$ 14.5% 25.6% 59.9% 4.9% 9.9% 13.6% 22.2% 49.4%
Table 3: Quantitative comparison of subject consistency, prompt following and visual quality.means publicly available version with Wan2.1-1.3B.
Methods Decoupled Concept Fidelity Prompt Video Quality
[…]
Vidu2.0 bao2024vidu 0.696 0.458 0.568 0.562 0.597 18.61 3.350 4.689
Pika2.1 pikaart2024 0.688 0.459 0.579 0.566 0.607 19.39 3.534 4.791
Kling1.6 kuaishou2024 0.659 0.420 0.552 0.547 0.582 18.38 3.487 4.787
Phantom liu2025phantom 0.703 0.476 0.589 0.573 0.615 17.73 3.404 4.812
Ours 0.744 0.533 0.598 0.600 0.644 18.87 3.565 4.903
Figure 4: Qualitative comparison with previous methods on subject consistency and text following.
Evaluation Metrics. As audio and appearance are key characteristics of humans, we follow current state-of-the-art method lin2025omnihuman ; huang2025conceptmaster to evaluate audio-driven, multi-concept human-centric video generation. To comprehensively evaluate methods for this task, we consider five dimensions: 1) Concept fidelity: We leverage CLIP-I radford2021learning and DINO-I oquab2023dinov2 on subjects in the generated videos to assess whether they are aligned with the provided reference images. To get cropped subject images in output videos, we randomly sample 5 frames in each video and use Florence-2 xiao2024florence to detect the subject bounding boxes. We then crop the detected regions and compute the CLIP-I and DINO-I scores. We also employ Face-Arc deng2019arcface , Face-Cur huang2020curricularface and Face-Glink deng2019arcface to evaluate the fidelity of facial features if the concept is a human. 2) Prompt following: We utilize video-level CLIP wang2022internvideo to measure the similarity between the input text prompts and the resulting video content. 3) Visual quality: We utilize q-align wu2023q , a vision-language model, for no-reference image quality assessment (IQA) and aesthetics score estimation (ASE). 4) Audio synchronization and human pose diversity: For lip synchronization, we leverage the widely used Sync-C and Sync-D syncnet to compute audio-visual confidence. We also incorporate hand keypoint confidence (HKC) and hand keypoint variance (HKV) lin2024cyberhost to quantify hand-pose accuracy and motion richness, respectively. For simplicity, we assume each test video contains exactly two participants: one speaking (with meaningful audio as input) and one listening (with muted audio as input). 5) Distribution distance: We employ FVD unterthiner2019fvd to measure the distance between generated and ground-truth videos.
Test Sets. We use three test sets in our experiments: 1) single-person audio conditioned human animation test set following OmniHuman lin2025omnihuman (see Tab. 1); 2) our collected two-person audio conditioned human animation test set where only one person is talking (see Tab. 1 and Tab. 4); and 3) multi-concept video customization test set following Phantom liu2025phantom , where we select 100 human-related pairs of reference images and text prompts in our experiments (see Tab. 3).
4.1 Comparisons with State-of-the-Arts
Multi-Concept Audio-Driven Human Video Generation. In Tab. 1, our approach achieves state-of-the-art or comparable performance in lip synchronization accuracy and motion diversity, particularly excelling in complex multi-person interactions where baseline methodslin2024cyberhost ; kuaishou2024 struggle with accurate audio signal assignments. For single-person scenarios, our method performs on par with specialized models like OmniHuman lin2025omnihuman . In multi-person settings, existing methods including OmniHuman and its extensions as well as leading commercial video generation models with post-processed lip-sync, fail to deliver satisfactory results. Poor lip-sync accuracy highlights their inability to generate accurate lip movements for the correct person. Although OmniHuman with oracle masks (manually setting audio-conditioned regions) improves lip-sync accuracy, it still significantly degrades overall video quality, as reflected in the FVD metric. This clearly demonstrates the limitations of existing methods that rely on single-identity assumptions, hindering their applicability to broader scenarios. In contrast, our method achieves strong performance in both lip-sync accuracy and overall video synthesis quality, effectively addressing the challenges of audio-conditioned multi-person video generation while remaining compatible with existing settings. Fig. 3 presents qualitative results for multi-person audio-driven video generation. While Kling1.6 w/ lip-sync shows many audio assignment errors and OmniHuman w/ fixed mask shows inflexible audio control with many missing cases, our method consistently assigns audio signals to the correct identity and demonstrates better motion dynamics and more precise audio-driven animation. It is worth noting that all previous methods rely on a reference frame containing complete information to generate talking-person videos, while our method only needs reference appearance of human’s head or full-body images and audios.
User Study. We conducted a user study to evaluate our method on two tasks: (1) lip synchronization in multi-person talking videos and (2) subject consistency in multi-concept customizations. The lip sync test used 19 videos from three methods, while subject consistency was assessed on 9 videos from five methods. We use the same model (labeled ‘ours’) for two evaluation tasks. Ten experienced users ranked each method with scores (higher values indicate superior performance). Tab. 2 reports the average scores and top-1 selection percentages. Our method achieved the highest scores and top-1 rates in both tasks, validating the effectiveness of our method.
Multi-Concept Video Customization. Although our major contribution is not on multi-concept video customization, we show that our model is also capable of preserving multi-concept visual appearances. In Tab. 3, our method outperforms existing approaches bao2024vidu ; pikaart2024 ; kuaishou2024 in preserving identity details and facial features, addressing the common degradation issue in multi-subject generation. Notably, we achieve this without sacrificing audio injection performance, while these methods are incapable to generate videos from audio conditions. It suggests our joint optimization framework on video generation and mask prediction successfully balances video generation. Our method ranks second in prompt following, likely due to our audio-driven human-centric training data focusing on talking and singing, which limits prompt diversity compared to text-to-video tasks. Despite this, qualitative results (Fig. 4) show that our approach maintains natural subject consistency and visual quality in both real-world and anime domains, outperforming previous methods that suffer from unnatural compositions and degraded visuals.
4.2 Ablation Study
We validate our local audio injection designs via ablation on three variants: 1) Global audio lin2025omnihuman applies audio across the entire feature map; 2) ID Embedding injects audio features with a learnable ID embedding without mask prediction; 3) Fixed Mask uses predefined static spatial masks. As shown in Tab. 4, our framework with predicted dynamic masks achieves the best Sync-D and FVD scores. The fixed mask yields decent Sync-D but suffers motion artifacts (worst FVD), and global audio, while scoring best in IQA, offers poor audio-visual alignment. Qualitative results in Fig. 5 further reveal that global audio drives all identities uniformly, ID embedding often mismatches audio with identities, and fixed masks lose alignment when characters move. These findings underscore that current methods overlook the need for precise local conditions, highlighting the strength of our dynamic, adaptive mask prediction strategy for multi-person talking video generation.
Table 4: Ablation study on audio-driven multi-person animation methods.
[…]
Global audio condition 9.482 4.768 3.466 33.895
ID Embedding 8.627 4.658 3.338 35.665
Fixed Mask 7.068 4.690 3.369 40.239
Predicted Mask (Ours) 6.670 4.757 3.467 22.881
Figure 5: Qualitative ablation on audio injection strategies.
5 Conclusion
In this paper, we introduced InterActHuman, a novel end-to-end video diffusion framework that supports spatially aligned, multi-modal conditioning for multi-concept human animation. By integrating an efficient mask-prediction module into a pretrained DiT backbone, our method automatically infers per-identity spatio-temporal layouts from reference images and uses these masks to guide local audio conditioning. This explicit layout binding enables each reference concept to retain its unique appearance and voice, even in complex scenes with multiple interacting entities. To facilitate training procedure of our model, we presented a high-throughput pipeline for harvesting and annotating over 2.6 million identity-aware video snippets, complete with per-frame masks for humans and objects. Extensive experiments on both single- and multi-person benchmarks demonstrate that InterActHuman achieves state-of-the-art performance in lip synchronization, motion diversity and subject apperance fidelity, while maintaining competitive video quality. We hope it could serve as a solid baseline for the multi-concept human animation and audio-driven multi-person video generation community, where further improvements could be built upon it.
Limitations. Since the goal of this work is for human-centric video generation, the available data domain is inherently narrower than one used for general text-to-video pretraining, limiting the ability to follow the diverse text prompts. While our framework is designed to accommodate any number of concept images, the training dataset predominantly consists of videos featuring two to three individuals. This distribution may constrain the generalization of any number of inputs, suggesting further improvements could be achieved by enlarging the diversity and scale of training data.
Broader Impacts. Our model could be used to generate misinformation by using celebrity images and voices. We will strictly restrict access and add watermarks to prevent misuse.
Appendix A Algorithm of Our Model Implementation
Please refer to Algorithm 1 for our model implementation.
Algorithm 1 InterActHuman: Audio-Driven Multi-Concept Video Generation Inference
1:
2:Text prompt $T$, reference images $\{X_{i}\}_{i=1}^{N}$, identity-level audio $\{Y_{i}\}_{i=1}^{N}$
3:Total diffusion steps $S$, mask injection threshold step $S_{\text{mask}}$
4:VAE Encoder/Decoder, diffusion transformer
5:Generated video $V$
6:Preprocessing:
7:for $i\leftarrow 1$ to $N$ do
8:     $T_{i}\leftarrow\text{Rephrase}(X_{i},T)$triangleright Generate detailed prompt via Qwen2.5-VL.
9:     $\mathbf{x}_{i}\leftarrow\text{VAE\_Encoder}(X_{i})$triangleright Encode reference image.
10:     $\mathbf{a}_{i}\leftarrow\text{wav2vec}(Y_{i})$triangleright Extract audio features.
11:end for
12:$c_{\text{text}}\leftarrow\{T,T_{1},\ldots,T_{N}\}$triangleright Aggregate text conditions.
13:$z_{S}\sim\mathcal{N}(0,I)$triangleright Initialize the noisy video latent.
14:Initialize mask cache: $\{m_{i}^{\text{prev}}\}\leftarrow 0$
15:for $k\leftarrow S$ downto $1$ do
16:     for each DiT block layer $l$ that we inject conditions do
17:         for $i\leftarrow 1$ to $N$ do
18:              Reference Injection: Inject $\{\mathbf{x}_{i}\}$ via concatnation and self-attention.
19:              Compute hidden video feature $\mathbf{h}^{v}$ and reference feature $\mathbf{h}^{r}_{i}$.
20:              $\mathbf{Q}^{v}\leftarrow\text{Proj}_{q}(\mathbf{h}^{v})$.  $[\mathbf{K}^{r}_{i},\,\mathbf{V}^{r}_{i}]\leftarrow\text{Proj}_{k,v}(\mathbf{h}^{r}_{i})$.
21:              Apply LayerNorm and 3D RoPE to the features.
22:              Compute cross-attention:
$\mathbf{p}^{(l)}_{i}\leftarrow\operatorname{softmax}\Big{(}\frac{\mathbf{Q}^{v}(\mathbf{K}^{r}_{i})^{\top}}{\sqrt{d}}\Big{)}\mathbf{V}^{r}_{i}.$
23:              Predict layer mask:
$m^{(l)}_{i}\leftarrow\operatorname{sigmoid}\Big{(}\text{MLP}(\mathbf{p}^{(l)}_{i})\Big{)}.$
24:         end for
25:     end for
26:     Aggregate Masks:
$m_{i}\leftarrow\frac{1}{L}\sum_{l=1}^{L}m^{(l)}_{i}\quad\forall~{}i\in\{1,\ldots,N\}.$
27:     Cache current masks: $\{m_{i}^{\text{prev}}\}\leftarrow\{m_{i}\}$.
28:     if $k<S_{\text{mask}}$ then
29:         for $i\leftarrow 1$ to $N$ do
30:              Audio Injection: inject local audio by updating the noisy latent $\mathbf{h}^{v}$.
31:              $\mathbf{Q}^{v}\leftarrow\text{Proj}_{q}(\mathbf{h}^{v})$.  $[\mathbf{K}^{\text{mute}}_{i},\,\mathbf{V}^{\text{mute}}_{i}]\leftarrow\text{Proj}_{k,v}(\mathbf{a}_{i}^{\text{mute}})$.  $[\mathbf{K}_{i},\,\mathbf{V}_{i}]\leftarrow\text{Proj}_{k,v}(\mathbf{a}_{i})$.
32:              Compute cross-attention:
$\mathbf{p}_{i}\leftarrow\operatorname{softmax}\Big{(}\frac{\mathbf{Q}^{v}(\mathbf{K}_{i})^{\top}}{\sqrt{d}}\Big{)}\mathbf{V}_{i},\quad\quad\mathbf{p}^{\text{mute}}_{i}\leftarrow\operatorname{softmax}\Big{(}\frac{\mathbf{Q}^{v}(\mathbf{K}^{\text{mute}}_{i})^{\top}}{\sqrt{d}}\Big{)}\mathbf{V}^{\text{mute}}_{i},$
33:              Bind audio conditions:
$\mathbf{h}^{v}\leftarrow\mathbf{h}^{v}+m_{i}\odot\mathbf{p}_{i}+(1-m_{i})\odot\mathbf{p}_{i}^{\text{mute}},$
where $\odot$ denotes element-wise multiplication.
34:         end for
35:     end if
36:     Update latent via diffusion step via flow-matching formulas or custom samplers.
37:end for
38:Decoding: Decode the final latent via VAE:
$V\leftarrow\text{VAE\_Decoder}(z_{0}).$
return $V$
Appendix B Experiment Details of Our Model
B.1 Implementation Details
Our model was trained for 10,000 steps on 32 A800 GPUs with a learning rate of 3e-5. We adopted the PyTorch framework combined with Fully Sharded Data Parallel (FSDP) to finetune the DiT model across multiple GPUs. Specifically, the model was partitioned such that different GPUs handled distinct portions of the model’s parameters. We configured the effective batch size so that every node—comprising 8 GPUs—processed 2 videos simultaneously. With 32 GPUs in total (i.e., 4 nodes), this resulted in an overall effective batch size of 8 videos.
B.2 Ablation Details
In Tab. 4 and Fig. 5 of the paper, we validate our local audio injection designs via ablation on three variants: 1) Global audio lin2025omnihuman applies audio across the entire feature map; 2) ID Embedding injects audio features with a learnable ID embedding without mask prediction; 3) Fixed Mask uses predefined static spatial masks.
For global audio, it is identical to the pretrained audio-driven model lin2025omnihuman . For fixed mask, it is a manually input rectangle mask upon the audio cross-attention of pretrained audio-driven model lin2025omnihuman . It requires the generated human being static and it needs the user to ensure there is only one person inside a mask.
For ID embedding, here we provide some implementation details. Inspired by detection transformers carion2020end , we introduce a set of $N$ learnable ID embeddings $\mathbf{E}_{query}\in\mathbb{R}^{N\times C}$. These embeddings serve as identity tokens for the $N$ individuals. Specifically, we add the same learnable ID embedding $\mathbf{e}_{query}\in\mathbb{R}^{C}$ to a paired reference image and audio segment. In this way, we expect the model to implicitly match an individual (defined by a reference image) and audio segments in the output video to get synchronized lip movements. For the ID embedding dataset preparation, we incorporate multi-person videos into training. We use Active Speaker Detection (ASD) to extract talking persons’ identities as a bounding box with identity-number. Then we use the bounding box and SAM2 masks to match the identity-number of each timestep of audio based on their IoU. Finally we use paired audio with timestep and reference image to generate video, and use flow-matching loss only to supervised it. No mask information is provided to the model and no mask supervision adopted.
In the main paper, our experimental results in ablation study indicate that implicit matching of individuals and audio segments is worse than explicit matching with layout-aligned mask prediction and audio injection. It is worth noting that although our implementation shows this evidence, implicit matching is not necessarily worse than explicit matching. We show that a straight-forward implementation of implicit matching cannot generate satisfactory results, yet there could be better implementations to improve this result in the future.
Appendix C Details of Audio-driven Base Model’s Architecture and Training
This appendix provides a detailed description of the network structure and training specifics for the pretrained audio-driven single-person video generation model, OmniHuman lin2025omnihuman . As noted in our main paper, the text-to-video base model undergoes post-pretraining for audio conditioning, following the OmniHuman lin2025omnihuman methodology. Subsequently, our multi-concept framework is built and trained on this foundation.
C.1 Audio-driven Base Model
Our foundational model is composed of a Variational Autoencoder (VAE) and a Latent Diffusion Transformer (DiT). The VAE includes an encoder, which compresses raw pixel data into a compact latent representation, and a decoder, which reconstructs the original pixel inputs from these latent features. The VAE achieves compression ratios of (4, 8, 8) for the (t, h, w) dimensions, respectively. Both the encoder and decoder utilize a temporally causal convolutional architecture, facilitating image and video compression across both spatial and temporal domains within the joint latent space. The denoising latents possess 16 channels.
The DiT Blocks are based on the dual-stream Diffusion Transformer (DiT)esser2024scaling . This transformer processes video and text tokens through multiple self-attention layers and feedforward networks (FFNs) to learn representations that are both shared and modality-specific. SwiGLU is employed as the activation function to enhance nonlinear modeling capabilities. Additionally, AdaSinglechen2023pixart is used for efficient timestep modulation. Moreover, two-thirds of the FFN weights in the deeper layers share parameters, creating a hybrid-stream design that preserves model capacity while substantially improving parameter efficiency.
C.2 Mixed Training Strategy
We utilize a multi-condition training strategy based on the framework established in OmniHuman lin2025omnihuman . Our method employs a two-stage progressive training scheme to develop the base model’s capabilities. In the initial stage, the model is trained solely on text-to-video (T2V) data to build fundamental video generation abilities. The second stage introduces audio-synchronized datasets to expand the model’s functionality to include audio-driven generation and reference image injection. For this second stage, we follow two core principles from OmniHuman: 1) Tasks with stronger conditioning can leverage tasks with weaker conditioning and their associated data to broaden the effective training dataset; 2) Tasks with stronger conditioning should be assigned proportionally lower training ratios. Guided by these principles, we first train the reference image injection capability before introducing the audio-driven generation objective. This staged methodology facilitates efficient knowledge transfer while ensuring stable training dynamics throughout the multi-condition learning process.
Appendix D Additional Details for Audio-driven Base Model Dataset
D.1 Dataset Curation
We initially filter the T2V data using rules 1 through 6. Following this, we apply rule 7 to further refine the audio-driven data.
1. Video Clip. We begin by using PySceneDetect ( https://github.com/Breakthrough/PySceneDetect) to identify and trim shot transitions and fades within video clips. After this process, all clips are standardized to a duration of 5 to 30 seconds.
2. Human. Utilizing the annotated video captions, we implement a rule-based system to detect keywords such as "people", "human", "men", "women", "girl", and "boy". If any of these keywords are found, the video is categorized as human-related. This technique ensures the dataset is effectively filtered to comprise only videos pertinent to human activities and interactions.
3. Subtitles. We use PaddleOCR (  https://github.com/PaddlePaddle/PaddleOCR) to detect subtitles in the videos and remove clips where subtitles change. This step guarantees that the dataset emphasizes continuous and consistent visual content, reducing distractions from textual variations and improving the data quality for subsequent tasks.
4. Visual Quality. We utilize Q-align wu2023q to evaluate the visual quality of the videos, filtering out clips that do not meet a predefined threshold. This procedure ensures the dataset maintains a high standard of visual clarity, which is essential for accurate analysis and robust model performance. By eliminating low-quality segments, we enhance the overall dependability and utility of the dataset.
5. Aesthetics. We employ Q-align wu2023q to assess the aesthetic appeal of the videos and discard clips falling below a set threshold. Through aesthetic quality assessment, we can filter out videos containing post-production elements, thereby improving the quality of the training dataset. This measure ensures the dataset is composed of natural and unaltered video content, which better represents real-world scenarios and boosts the robustness and generalization of the trained models.
6. Motion. We use Raft teed2020raft to calculate the optical flow of the videos and filter out clips exhibiting overly intense motion. This step ensures that the dataset includes only video clips with moderate and significant motion, which are more appropriate for analysis and model training. By removing clips with extreme motion, we enhance the stability and quality of the dataset, leading to more dependable results.
7. Syncnet. For audio-driven content, we employ SyncNet chung2017out to determine if the lip movements are synchronized with the audio. Videos displaying considerable asynchronization are excluded. This step ensures that the dataset consists only of high-quality, synchronized audio-visual data. By removing out-of-sync segments, we improve the overall quality and reliability of the dataset. Ultimately, we gather 2,000 hours of audio-driven data.
D.2 Dataset Analysis
To gain a better understanding of the dataset’s distribution, we perform an analysis across three dimensions. Detailed definitions for each dimension are provided below.
  • Human Size. Human size indicates the portion of the human body visible within the video frame. It is categorized into the following levels: Portrait (head and shoulders), Chest (from head to chest), Waist (from head to waist), Knees (from head to knees), and Full Body (the entire body is visible). To determine this, we first detect body keypoints using RTMpose jiang2023rtmpose , and then classify the human size based on the confidence scores of these keypoints. This method ensures a robust and accurate categorization of the visible human body extent in each video.
  • Motion Amplitude. After extracting body keypoints from the video, the amplitude of human motion is computed by measuring the displacement of the chest keypoint over time, relative to the width of the shoulders. Based on these calculations, we classify motion amplitude into four categories: Slight($<0.1$), Moderate($0.1-0.2$), Significant($0.2-0.3$), and Extreme($>0.3$).
  • Scene. Using video captioning, we employ Doubao to categorize videos into the following scenarios: Household, Work, Show, Outdoor Adventure, transportation, Arts and Crafts, Sports, and Others.
  • Language. For data containing audio, we use a language detection tool ( https://github.com/Mimino666/langdetect) to identify and count the types of languages present. These are categorized as English, Chinese, Spanish, French, German, Urdu, Hindi, and Others.
Appendix E User Study Details
Here we show our questionnaires used in the user study.
E.1 Multi-person Audio-Driven Video Generation Questionnaire
Please take 10 minutes to complete the following ranking questions. For each question, consider the priority of the three videos based on the following dimensions and rank them accordingly:
(1) Lip-sync Accuracy: Based on the audio you hear, do you think the lip movements are accurate? Ideally, only one person’s lip movements should correspond to the audio, but it doesn’t matter specifically who is speaking. It is considered poor if multiple people are speaking simultaneously or if no one is speaking when there is audio. When multiple methods all correctly show only one person speaking, rank them based on the quality of lip-sync with the audio. Perfect synchronization is good; missing syllables or incorrect lip shapes for certain syllables is poor.
(2) Sense of Dialogue Among Multiple People: Is there a feeling of conversation between the individuals? It is considered good if it portrays a natural scenario where one person speaks and another listens. The judgment here is based on whether their expressions during the interaction appear natural and whether the overall feeling of the conversation is natural.
(3) Video Quality: If all the above criteria are tied, then prioritize the video with higher quality.
The priority of these three criteria decreases in the order listed. Aspect ratio and resolution should not be taken into account; you should not favor a particular aspect ratio or higher resolution, but rather focus on the inherent video quality itself.
E.2 Multi-concept Video Generation Questionnaire
Please take 8 minutes to complete the following ranking questions. For each question, consider the priority of the five videos from the perspective of consistency between the reference images and the appearance in the video, and rank them accordingly:
(1) Consistency with Reference Images: Based on the reference images, do you think all the listed reference images appear in the video? And does the appearance of those in the video match the images? Consider the priority in the following order from high to low:
Ranked highest: Appearance is consistent, and all reference images appear.
Next: All reference images appear, but the appearance of some reference images is inconsistent.
Ranked lowest: Some reference images do not appear.
(2) Video Quality: If the above criteria are tied, then prioritize the video with higher quality.
(3) Motion Strength: If all the above criteria are tied, then prioritize the video with a larger motion strength.
The priority of these criteria decreases in the order listed. Aspect ratio and resolution should not be taken into account; you should not favor a particular aspect ratio or higher resolution, but rather focus on the inherent video quality itself.